OpenStudy (anonymous):

Write a polynomial function in standard form with the given zeros x=2,-2,-3

OpenStudy (anonymous):

the factored form of the simplest polynomial with those zeros is: (x-2)(x+2)(x+3), all you have to do is expand and you get your answer

OpenStudy (anonymous):

i need to see how to work it out plz

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

they tell you that 2, -2, and -3 are zeros so you know that the factored form of the simplest polynomial with those zeros is (x-2)(x+2)(x+3) because x-2=0, x+2=0, x+3=0 expanding this yields (x^2-4)(x+3) -> x^3+3x^2-4x-12